# On the Challenges and Opportunities in Generative AI

### v3

@@ 4 Ethical Deployment and Societal Impact
With the current excitement around the scope and application of large-scale generative models, we are also witnessing a growing apprehension, fueled by media reports, of adverse outcomes surrounding the rapid advancement of generative AI. These concerns add to the conceptual and practical considerations discussed so far and encompass a range of issues, including the spread of misinformation, the absence of regulatory frameworks (Meskó & Topol, 2023), unintended harm (Greenfield & Bhavnani, 2023), and debates over open-source versus closed-source technologies (Chen et al., 2023), among others. Here we identify key challenges concerning the responsible deployment of large-scale deep generative models. More specifically, we discuss several aspects, including the dissemination of misinformation (4.1), violation of privacy and copyright (4.2), presence of biases (4.3), lack of interpretability (4.4), and constraint satisfaction (Section 4.5).
@@ 4.1 Misinformation and Uncertainty
As the quality of generated data synthesized using large-scale generative models increases, it can become more and more difficult to distinguish between real and generated content, especially for uninformed consumers (Frank et al., 2024). This indistinguishability facilitates the spread of misinformation (e.g., by deepfakes (Helmus, 2022)). To ensure the trustworthiness of information, we need algorithmic solutions that are on par with the advances in generative models and allow us to robustly detect and mark synthetic data. Watermarking is one such approach in which there has recently been increased interest. The goal of these methods is to manipulate a generated sample (e.g., an image or piece of text) such that a signature can be detected in downstream tasks albeit with minimal effects on sample quality. There have been several approaches to watermark synthetic data generated from LLMs (Kirchenbauer et al., 2023; Dathathri et al., 2024; Zhao et al., 2024) and image generation models (Zhao et al., 2023c; Wen et al., 2023; Jiang et al., 2024b). However, current watermarking methods are far from perfect (Saberi et al., 2023). Evasion is often possible by small manipulations (like paraphrasing or pixel perturbations) (Jiang et al., 2023) and the injecting of watermarks into content can be inefficient (Liu et al., 2024). While some research suggests that robust watermarking is possible (Krishna et al., 2023), more research is needed on the theoretical limitations of watermarking (Kirchenbauer et al., 2024).
@@
Notably, misinformation can emerge even without malicious intent. Tools like ChatGPT are increasingly expected to serve as universal question-answering engines, even though their core objective—to estimate the likelihood of the next token in a sequence—is traditionally designed to assess the linguistic plausibility of strings (Kalai & Vempala, 2023), rather than their factual accuracy. This distinction between the two objectives is evinced by the discrepancies observed between the probability a model explicitly assigns to a statement when prompted vs. the underlying log-probability it assigns (Hu & Levy, 2023), models’ tendencies to hallucinate (Huang et al., 2024a)), and their difficulty in achieving probabilistic consistency (Elazar et al., 2021), e.g., ensuring logical predictions between a statement and its negation.
@@
Some works have turned to model uncertainty estimates (e.g., those discussed in 3.2) as indicators of model reliability, developing methods to enhance the trustworthiness of AI systems based on these estimates (Edupuganti et al., 2021; Yang et al., 2023c). For example, (Ren et al., 2023) propose a selective generation approach, where models abstain from providing a response in the face of high uncertainty; (Kuhn et al., 2023) use a notion of a model’s semantic uncertainty to predict the correctness of its answer in question-answering. The development of methods that explicitly account for uncertainty represents an interpretable approach toward ensuring model reliability, offering a strategy that also has grounding in a well-studied concept in machine learning (Malinin & Gales, 2018; Abdar et al., 2021; Gawlikowski et al., 2023).
@@
Encouragingly, some research suggests that larger LMs actually are well-calibrated in terms of their world knowledge, i.e., their predicted likelihoods reflect the probability that a statement is true (Srivastava et al., 2022; Zhu et al., 2023; Yu et al., 2024). Further, recent studies show that language models often do possess the ability to assess the truthfulness of their own statements (Lin et al., 2022; Kadavath et al., 2022; Xiong et al., 2024). However, fine-tuning or RLHF, which are frequently applied to these models, have been shown to hurt calibration; rather, they have been widely observed to exhibit overconfidence—the tendency of a model to assign excessively high probabilities to its predictions regardless of their correctness (Kadavath et al., 2022; Tian et al., 2023; OpenAI, 2023; Xiong et al., 2024). There has been some work on mitigating miscalibration issues for fine-tuning (e.g., Wang et al., 2023b) and RLHF (e.g., Tian et al., 2023; Zhang et al., 2024). and there is an increasing focus on systems where evidence can be brought in from external knowledge sources (Blattmann et al., 2022; Pan et al., 2023; Gao et al., 2023), grounding model responses to reliable knowledge sources. Such research—along with benchmarks to assess model factuality (e.g., KoLA; Yu et al., 2024)—is a critical step towards ensuring the trustworthiness and reliability of generative models.
4.2 Privacy and Copyright Infringement
@@
Interestingly, recent works show that publicly available LLMs and large-scale text-to-image models can implicitly “memorize” training data (van den Burg & Williams, 2021), to the point that samples from the dataset can be (almost exactly) reconstructed (Carlini et al., 2023a; b; Somepalli et al., 2023; Nasr et al., 2023). This behavior potentially infringes on data privacy, underscoring the importance of exploring whether generative models can be trained while safeguarding sensitive information. Approaches for alleviating this problem have been proposed: for example, differential privacy (DP) constraints, which can be enforced during generative model training, offer an attractive theoretical framework to ensure privacy Li et al. (2021); Dockhorn et al. (2023). However, DP-based approaches suffer from a tradeoff between privacy and utility (Cummings et al., 2024). Moreover, in the context of image generation, scaling such approaches to high-resolution datasets remains elusive. Recent work has instead focused on generative DP synthetic data, where foundation models are only used as blackboxes (Lin et al., 2024b). Therefore, building privacy constraints in large-scale training of generative models can be a promising direction for further research.
Another byproduct of memorization in generative models is that it can lead to unauthorized distribution or replication of training data, resulting in copyright infringement liabilities. [^8] https://www.nytimes.com/2023/12/27/business/media/new-york-times-open-ai-microsoft-lawsuit.html While some recent work tries to alleviate this issue (Liang et al., 2023; Zhao et al., 2023d), there are several outstanding technical challenges ranging from mitigating copyright infringements during dataset curation (Carr & Jeffrey, 2022; Duarte et al., 2024) to reliable detection of copyright violations in generated samples (Li et al., 2024a; Kim et al., 2021). Beyond privacy and copyright concerns, memorization has other undesirable effects. Many of the use cases that generative models are employed for are dynamic in nature, e.g., creative writing or graphics generation; outputs that simply replicate training data lack diversity and originality, and may fail to address user-specific needs or requests. This behavior potentially creates a worse user experience and limits models’ usefulness. Further, excessive memorization can lead to biases in generated content, where certain perspectives or styles dominate because they were overrepresented in the training data; we discuss this last issue in more detail in the next section.
@@ 4.3 Fairness
Not all unfair behaviors exhibited by generative models are as overtly harmful as generating toxic content. Models can show more subtle biases towards certain subpopulations, such as allocation bias—when AI systems extend or withhold opportunities, resources, or information—or quality-of-service bias—when AI systems work better for people in some subpopulations than others; these behaviors should not be downplayed as they can perpetuate systemic inequalities. A main culprit of such behaviors stems from the fact that the data used in the training of most generative models is disproportionately from certain countries and languages. Such models therefore might not work as well for languages or images outside of these mainstream groups. For example, multilingual language models work substantially better for English use cases (Lai et al., 2023; Huang et al., 2023a). Further, because the tokenizers for such models have also been trained disproportionately using English data, the compression rate for English texts is much higher than for texts in low-resource languages (Petrov et al., 2023; Ahia et al., 2023). Consequently, services that charge based on token counts impose higher costs for a query made in a low-resource language compared to a query with the same underlying meaning made in English. This exacerbates accessibility challenges for speakers of underrepresented languages.
@@
While numerous approaches have been proposed to mitigate the biases of generative models in a post-hoc manner (Bai et al., 2022; Glaese et al., 2022; Ferrara, 2024; Olmos et al., 2024), the achieved changes are often merely superficial, leaving the possibility of remnant biases. For example, Gonen & Goldberg (2019) demonstrate that word embeddings still cluster based on gender stereotypes even after bias mitigation techniques, effectively “hiding” rather than eliminating the issue. In the vision domain, post-hoc bias mitigation strategies have been observed to work poorly in the face of test-time distribution shift (Kong et al., 2023). Further, the efficacy of these methods has in most cases only been verified for a small subset of bias or fairness criteria. More research is necessary both to train models that can readily account for multiple forms of fairness with regard to diverse criteria and to fully understand the strengths and shortcomings of bias mitigation strategies for generative AI.
@@
Ultimately, assessing and ensuring fairness in technology applications is a complex challenge. Aside from the aforementioned issue of differing (and potentially dynamic) qualitative definitions of fairness (3.2), different notions of fairness often cannot be fully satisfied simultaneously (Ferrara, 2024). Therefore, it is essential for the builders of generative AI tools to carefully evaluate the various dimensions of fairness and make deliberate trade-offs appropriate for the specific use .
@@ 4.4 Interpretability and Transparency
In many situations, it is critical to understand the logic and influencing factors behind generative models’ outputs. In other words, we need to be able to interpret how a generative model produces its results, with its decision-making process being transparent, i.e., accessible and understandable This is particularly true in safety-critical domains—such as healthcare (Chen et al., 2021a) or finance applications—but is also important across general AI use cases, where interpretability is essential for diagnosing errors and fostering user trust. These needs are not new and have been present since the start of publicly-available AI-based products and tools (Confalonieri et al., 2021). There has thus been a sizable amount of research in neural network interpretability methods. However, these methods are not always feasible for use with large-scale DGMs. For example, interpretability methods, such as SHAP, LIME, or Integrated Gradients (Lundberg & Lee, 2017; Ribeiro et al., 2016; Sundararajan et al., 2017), struggle to scale effectively with the complexity and size of large models; many interpretability methods work by attempting to understand concepts encoded in models’ latent representations (Crabbe et al., 2021; Esser et al., 2020), but this becomes more difficult in the high-dimensional latent spaces used by DGMs. Further, while one might hope that explanations derived from interpreting smaller models could be used for understanding their larger counterparts, scaling up generative models gives rise to unpredictable effects, e.g., models demonstrating unexpectedly advanced capabilities (Wei et al., 2022; Schaeffer et al., 2024); conclusions drawn when using small models therefore may not be applicable to today’s larger models.
@@
The fundamental challenge is to develop explanation methods for DGMs that are both well-understood by humans and faithful to the underlying model behaviors (Schut et al., 2023; Gurnee & Tegmark, 2024). Mechanistic interpretability is an emerging field that attempts to achieve this goal by reverse-engineering neural network decisions, translating them to human-interpretable decision-making processes (Bereska & Gavves, 2024). This is specifically done by analyzing models at the level of their internal computations, representations, and structural components. A large appeal of mechanistic interpretability is that provides causal explanations for models’ outputs, i.e., a decision or prediction can be attributed in a causal manner to some component of the input. This allows researchers to move beyond correlation-based explanations and instead understand the actual computational mechanisms driving observed model behavior, which in turn enables more precise debugging, bias detection, and control over generative outputs. Mechanistic interpretability research has uncovered a number of interesting and useful properties of DGMs. For example, specific neurons or layers in GANs and VAEs encode “disentangled” (i.e., distinct and interpretable) features, such as shape, texture, or pose in images (Shen et al., 2020; Mita et al., 2021). Other works have found that certain attention heads in transformers correspond to meaningful linguistic patterns, e.g., some might focus on syntactic structure while others might capture semantic information (Vig & Belinkov, 2019; Elhage et al., 2021). Such properties not only help practitioners better understand DGMs, they also enable them to control generation to some extent (Härkönen et al., 2020; Fetty et al., 2020).
@@
However, the reliability and comprehensiveness of mechanistic interpretability methods remain a subject of debate. A central criticism is that, while these techniques aim to identify causal relationships between model components and outputs, they may only provide partial or even misleading insights into the actual computational processes at work. For instance, attention patterns—which have been used by various methods to attribute model predictions to certain tokens (Xu et al., 2015; Choi et al., 2016, inter alia)—do not always faithfully reflect how or why certain tokens influence the final prediction (Jain & Wallace, 2019; Liu et al., 2022b). In some cases, they may merely highlight correlations rather than reveal deeper causal structures. Similar concerns have been raised about other methods for explaining model behaviors from mechanistic interpretability. Further, the new wave of large-scale generative models makes the application of some of these methods more difficult: the larger a model becomes, the (arguably) more difficult it becomes to fully reverse-engineer a prediction. Every nuance of these models’ decision-making may not be deducible from a subset of neurons, layers, or attention heads; interpretations derived from one subset of model components may overlook equally critical interactions elsewhere in the large network. Representation engineering (Zou et al., 2023a) and the use of sparse autoencoders (Bricken et al., 2023) are lines of research in mechanistic interpretability that potentially address these issues, offering interpretable explanations even for large-scale models. However, we are still in need of validation methods that can confirm whether the identified “mechanisms” truly govern a model’s outputs, or whether they merely reflect convenient, yet incomplete, narratives about its internal workings.

### v4

@@ 4 Ethical Deployment and Societal Impact
Sub-challenge Typical failure mode Promising mitigation avenues Reference materials Research-question candidates
Misinformation & synthetic media detection (§4.1) Deepfakes bypass detectors Model-rooted watermarks (e.g., Tree-Ring Watermarks) Datasets:
Deepfake Detection Challenge
; WaterBench
Surveys:
Rana et al. (2022)
Q1: Can diffusion-time watermarking survive multimodal adversarial attacks?
Q2: Can uncertainty-aware abstention policies mitigate hallucinated facts in DGM outputs?
Privacy violations, copyright infringement (§4.2) & PII leakage Model regenerates training data verbatim or near-verbatim Differential Privacy learning techniques (e.g., DP-SGD); Machine Unlearning methods Datasets:
CPDM
; PrivLM-Bench
Surveys:
Yao et al. (2024)
Kibriya et al. (2024)
Q1: What privacy-preserving fine-tuning strategies remain feasible for $\geq 100$B-parameter models under practical compute budgets?
Fairness across languages (§4.3) Worse compression $\rightarrow$ more tokens needed $\rightarrow$ higher cost for low-resource languages Dynamic tokenization schemes; Vocabulary transfer methods Surveys:
Xu et al. (2025)
Qin et al. (2025)
Q1: Can subword-free sequence modeling architectures (e.g., SSMs) eliminate tokenization-induced performance disparities across languages?
Bias & discrimination in generated content (§4.3) Models reflect and propagate bias and discrimination present in training data Counterfactual evaluation benchmarks; Fairness metric (e.g., demographic parity) incorporation into training objectives Datasets:
HolisticBias
; OpenBias
Surveys:
Gallegos et al. (2024)
Q1: How do watermarking methods interact with the appearance of demographic biases in model outputs?
Interpretability & transparency (§4.4) DGM parameters are uninterpretable, making them black boxes for users and developers Automated circuit discovery; Causal tracing Datasets:
GPT-2 neuron-explanation dataset
Surveys:
Marcinkevics & Vogt (2020)
Q1: Which confidence metrics best predict whether a mechanistic explanation truly modulates model behavior?
Q2: Can mechanistic insights be transferred across model sizes?
Constraint satisfaction (§4.5) Outputs violate hard rules (e.g. code won’t compile) Context Free Grammar-guided decoding algorithms; Constrained RLHF; Constraint-embedded model architectures Benchmarks:
HumanEval
; BigCodeBench
Surveys:
Zhang et al. (2023a)
Q1: How can hard-constraint decoding be generalised from code to text and images?
Q2: Do constraint-aware training methods (e.g., constrained RLHF) achieve better safety-performance trade-offs than inference-time constraint enforcement (i.e., constrained decoding)?
Table 3: Research challenges and mitigation strategies - Ethical Deployment and Societal Impact
With the current excitement around the scope and application of large-scale generative models, we are also witnessing a growing apprehension, fueled by media reports, of adverse outcomes surrounding the rapid advancement of generative AI. These concerns add to the conceptual and practical considerations discussed so far and encompass a range of issues, including the spread of misinformation, the absence of regulatory frameworks (Meskó & Topol, 2023), unintended harm (Greenfield & Bhavnani, 2023), and debates over open-source versus closed-source technologies (Chen et al., 2023), among others. Here we identify key challenges concerning the responsible deployment of large-scale DGMs. More specifically, we discuss several aspects, including the dissemination of misinformation (4.1), violation of privacy and copyright (4.2), presence of biases (4.3), lack of interpretability (4.4), and constraint satisfaction (Section 4.5).
@@ 4.1 Misinformation and Uncertainty
As the quality of generated data synthesized using large-scale generative models increases, it can become more and more difficult to distinguish between real and generated content, especially for uninformed consumers (Frank et al., 2024). This indistinguishability facilitates the spread of misinformation (e.g., by deepfakes (Helmus, 2022)). To ensure the trustworthiness of information, we need algorithmic solutions that are on par with the advances in generative models and allow us to robustly detect and mark synthetic data. Numerous models for differentiating machine-generated from real content have been proposed over the last years (Rana et al., 2022), but the increasing quality of generative model outputs has decreased their accuracy. Watermarking is another approach in which there has recently been increased interest. The goal of these methods is to manipulate a generated sample (e.g., an image or piece of text) such that a signature can be detected in downstream tasks, albeit with minimal effects on sample quality. There have been several approaches to watermark synthetic data generated from LLMs (Kirchenbauer et al., 2023; Dathathri et al., 2024; Zhao et al., 2024) and image generation models (Zhao et al., 2023c; Wen et al., 2023; Jiang et al., 2024b). However, current watermarking methods are far from robust (Saberi et al., 2023). Evasion is often possible by small manipulations (like paraphrasing or pixel perturbations) (Jiang et al., 2023) and the injecting of watermarks into content can be inefficient (Liu et al., 2024a). Recent work has demonstrated that model-integrated watermarks—i.e., signatures embedded in a model’s sampling process rather than post-hoc in its outputs—are a promising route forward, as they can survive a range of real-world corruptions (e.g., paraphrasing attacks in text (Krishna et al., 2023) and latent-trajectory perturbations in images (Wen et al., 2023; Fernandez et al., 2023)). Nonetheless, the information-theoretic limits on detectability, false-positive rate, and adversarial removability for watermarking remain poorly understood, and initial negative results suggest unavoidable trade-offs between robustness, the quality of the altered sample, and watermark payload (Kirchenbauer et al., 2024; Yoo et al., 2024).
@@
Notably, misinformation can emerge even without malicious intent. Tools like ChatGPT are increasingly expected to serve as universal question-answering engines, even though their core objective—to estimate the likelihood of the next token in a sequence—is traditionally designed to assess the linguistic plausibility of strings (Kalai & Vempala, 2023), rather than their factual accuracy. This distinction between the two objectives is evinced by the discrepancies observed between the probability a model explicitly assigns to a statement when prompted vs. the underlying log-probability it assigns (Hu & Levy, 2023), models’ tendencies to hallucinate (Huang et al., 2024a), and their difficulty in achieving probabilistic consistency (Elazar et al., 2021), e.g., ensuring logical predictions between a statement and its negation.
@@
Some works have turned to model uncertainty estimates (e.g., those discussed in 3.2) as indicators of model reliability, developing methods to enhance the trustworthiness of AI systems based on these estimates (Edupuganti et al., 2021; Yang et al., 2023e). For example, Ren et al. (2023) propose a selective generation approach, where models abstain from providing a response in the face of high uncertainty; Kuhn et al. (2023) use a notion of a model’s semantic uncertainty to predict the correctness of its answer in question-answering. The development of methods that explicitly account for uncertainty represents an interpretable approach toward ensuring model reliability, offering a strategy that also has grounding in a well-studied concept in machine learning (Malinin & Gales, 2018; Abdar et al., 2021; Gawlikowski et al., 2023).
@@
Encouragingly, some research suggests that larger LMs actually are well-calibrated in terms of their world knowledge, i.e., their predicted likelihoods reflect the probability that a statement is true (Srivastava et al., 2022; Zhu et al., 2023; Yu et al., 2024). Further, recent studies show that language models often do possess the ability to assess the truthfulness of their own statements (Lin et al., 2022; Kadavath et al., 2022; Xiong et al., 2024). However, fine-tuning or RLHF, which are frequently applied to these models, have been shown to hurt calibration; rather, they have been widely observed to exhibit overconfidence—the tendency of a model to assign excessively high probabilities to its predictions regardless of their correctness (Kadavath et al., 2022; Tian et al., 2023; OpenAI, 2023; Xiong et al., 2024). There has been some work on mitigating miscalibration issues for fine-tuning (e.g., Wang et al., 2023b) and RLHF (e.g., Tian et al., 2023; Zhang et al., 2024), and there is an increasing focus on systems where evidence can be brought in from external knowledge sources (Blattmann et al., 2022; Pan et al., 2023; Gao et al., 2023), grounding model responses to reliable knowledge sources. Such research—along with benchmarks to assess model factuality (e.g., KoLA; Yu et al., 2024)—is a critical step towards ensuring the trustworthiness and reliability of generative models.
4.2 Security, Privacy and Copyright Infringement
While modern generative models like LLMs are deployed practically for many applications, this also exposes them to potential malicious attacks, which can have significant costs for downstream applications or users. One class of malicious attacks on LLMs is the so-called “backdoor attacks” where the main idea is to train the model using poisoned data and then trigger a specific output response from the model corresponding to specific prompts. For instance, Yang et al. (2023a) discusses backdoor attacks on LLMs in the context of communication networks (see Zhou et al., 2025, for a more in-depth exploration of backdoor attacks). Another class of attacks known as “jail-breaking” involves designing adversarial prompts to generate malicious outputs from the model while bypassing guardrails employed to comply with usage policies. There has been a good deal of research in the context of LLMs exploring techniques for jail-breaking and for guarding against jail-breaking (Shen et al., 2024; Yi et al., 2024; Jin et al., 2024).
@@ 4.2 Security, Privacy and Copyright Infringement
Beyond malicious attacks, another critical risk associated with generative AI models is their tendency to memorize and unintentionally reproduce training data, raising serious concerns about privacy and data leakage. A variety of works have shown that publicly available LLMs and large-scale text-to-image models can implicitly “memorize” training data (van den Burg & Williams, 2021), to the point that samples from the dataset can be (almost exactly) reconstructed (Carlini et al., 2023a; b; Somepalli et al., 2023; Nasr et al., 2023). This behavior potentially infringes on data privacy, underscoring the importance of detecting whether private information has been leaked into an LLM’s training data (Kim et al., 2023) and exploring whether generative models can be trained while safeguarding sensitive information. Differential privacy (DP) constraints, which can be enforced during generative model training, offer an attractive theoretical framework to ensure privacy Dwork & Roth (2014); Li et al. (2021); Dockhorn et al. (2023). However, DP-based approaches have several shortcomings. They suffer from a trade-off between privacy and utility (Cummings et al., 2024). There are different DP formulations, each based on different assumptions about trust, data access, and the point at which noise is introduced. The meanings of the canonical DP parameters are thus not consistent, making comparison of models produced using different approaches difficult (Li et al., 2024d). Recent work has instead focused on generative DP synthetic data, where foundation models are only used as black boxes (Lin et al., 2024b). Building privacy constraints into the training of large-scale generative models can be a promising direction for further research.
Another byproduct of memorization in generative models is that it can lead to unauthorized distribution or replication of training data, resulting in copyright infringement liabilities. [^9] https://www.nytimes.com/2023/12/27/business/media/new-york-times-open-ai-microsoft-lawsuit.html Current efforts towards preventing such behavior have approached the problem from different angles. Some focus on filtering training data: techniques such as contractual licensing filters and hash-based de-duplication allow developers to identify and exclude protected material before training (Carr & Jeffrey, 2022; Duarte et al., 2024). Plagiarism or style-clone detectors allow intervention downstream, flagging generated samples that are substantially similar to copyrighted works (Li et al., 2024c; Kim et al., 2021). Other works have proposed training methods, which can be broadly categorized under two complementary strategies: (i) imitation-resistant training objectives that discourage verbatim memorization (Liang et al., 2023; Zhao et al., 2023d), and (ii) machine unlearning methods that aim to eliminate the influence of certain datapoints (e.g., copyrighted or private material) from a model’s predictions after training (Li et al., 2024b; Liu et al., 2024b; 2025). There is still much open research in this domain, including provenance tracking that can scale to work with modern, massive datasets and copyright infringement risk metrics that encompass legal concerns.
Beyond privacy and copyright concerns, indiscriminate memorization has other undesirable effects. Many of the current target applications for generative models—such as creative writing or graphics generation—demand novelty and user-specific adaptation; when a model merely regurgitates training data, it does not provide the desired diversity, originality and personalization. This behavior potentially worsens user experience and limits the practical value of generative models. Further, excessive memorization can lead to biases in generated content, where certain perspectives or styles dominate because they were overrepresented in the training data; we discuss this last issue in more detail in the next section. Ultimately, deploying DGMs safely and usefully will require turning the request of “don’t copy the training set” into an explicit design and auditing objective, which necessitates continued research into how to measure, control, and (when necessary) unlearn memorized examples (Chen et al., 2024a; Lesci et al., 2024).
@@ 4.3 Fairness
Not all unfair behaviors exhibited by generative models are as overtly harmful as generating toxic content. Models can show more subtle biases towards certain subpopulations, such as allocation bias—when AI systems extend or withhold opportunities, resources, or information—or quality-of-service bias—when AI systems work better for people in some subpopulations than others; these behaviors should not be downplayed as they can perpetuate systemic inequalities. A main culprit of such behaviors stems from the fact that the data used in the training of most generative models is disproportionately from certain countries and languages. Such models therefore might not work as well for languages or images outside of these mainstream groups. For example, multilingual language models typically perform substantially better on English tasks compared to tasks in other languages (Lai et al., 2023; Huang et al., 2023a). Further, because the tokenizers for such models have also been trained disproportionately using English data, the compression rate for English texts is much higher than for texts in low-resource languages (Petrov et al., 2023; Ahia et al., 2023). Consequently, services that charge based on token counts impose higher costs for a query made in a low-resource language compared to a query with the same underlying meaning made in English. This exacerbates accessibility challenges for speakers of underrepresented languages.
@@
Numerous approaches have been proposed to mitigate the biases of generative models in a post-hoc manner (Bai et al., 2022; Glaese et al., 2022; Ferrara, 2024; Olmos et al., 2024). However, the achieved changes are often merely superficial, leaving the possibility of remnant biases. For example, Gonen & Goldberg (2019) demonstrate that word embeddings still cluster based on gender stereotypes even after bias mitigation techniques, effectively “hiding” rather than eliminating the issue. In the vision domain, post-hoc bias mitigation strategies have been observed to work poorly in the face of test-time distribution shift (Kong et al., 2023). Moreover, most evaluations assess only a single fairness axis—for instance, gender in English or skin-tone in photos—and residual biases along other dimensions can remain undetected. Some key research areas that we identify as needing further attention include: (i) joint evaluation across multiple, potentially interacting fairness criteria (e.g., gender $\times$ dialect); (ii) stress-testing mitigation strategies under data-distribution shifts; and (iii) training-time interventions that prevent harmful biases from emerging in the first place. Some areas that we identify as needing further research are combined evaluation with respect to multiple forms of fairness criteria, robust assessments across multiple domains and training methods that can more robustly mitigate the learning of harmful biases. Promising steps in these directions include multilingual, multi-attribute benchmark suites such as XCOPA-Bias (Goldfarb-Tarrant et al., 2023) and gradient-based de-biasing schedules that adjust sampling weights during pre-training (Kim et al., 2024).
@@
Ultimately, assessing and ensuring fairness in technology applications is a complex challenge. Aside from the aforementioned issue of differing (and potentially dynamic) qualitative definitions of fairness (3.2), different notions of fairness often cannot be fully satisfied simultaneously (Ferrara, 2024). Therefore, it is essential for the builders of generative AI tools to carefully evaluate the various dimensions of fairness and make deliberate trade-offs appropriate for the specific use case.
@@ 4.4 Interpretability and Transparency
In high-stakes applications such as healthcare and legal domains, it is critical to understand the logic and influencing factors behind generative models’ outputs. In other words, we need to be able to interpret how a generative model produces its results, with its decision-making process being transparent, i.e., accessible and understandable This is particularly true in safety-critical domains—such as healthcare (Chen et al., 2021a) or finance applications—but is also important across general AI use cases, where interpretability is essential for diagnosing errors and fostering user trust. These needs are not new and have been present since the start of publicly available AI-based products and tools (Confalonieri et al., 2021). There has thus been a sizable amount of research in neural network interpretability methods. However, these methods are not always feasible for use with large-scale DGMs. For example, interpretability methods, such as SHAP, LIME, or Integrated Gradients (Lundberg & Lee, 2017; Ribeiro et al., 2016; Sundararajan et al., 2017), struggle to scale effectively with the complexity and size of large models; many interpretability methods work by attempting to understand concepts encoded in models’ latent representations (Crabbe et al., 2021; Esser et al., 2020), but this becomes more difficult in the high-dimensional latent spaces used by DGMs. Further, while one might hope that explanations derived from interpreting smaller models could be used for understanding their larger counterparts, scaling up generative models gives rise to unpredictable effects, e.g., models demonstrating unexpectedly advanced capabilities (Wei et al., 2022; Schaeffer et al., 2024); conclusions drawn when using small models therefore may not apply to today’s larger models.
@@
The fundamental challenge is to develop explanation methods for DGMs that are both well-understood by humans and faithful to the underlying model behaviors (Schut et al., 2023; Gurnee & Tegmark, 2024). Mechanistic interpretability is a field that attempts to achieve this goal by reverse-engineering neural network decisions, translating them to human-interpretable decision-making processes (Bereska & Gavves, 2024). This is specifically done by analyzing models at the level of their internal computations, representations, and structural components, e.g., identifying minimal subnetworks (referred to as circuits) that implement a specific computation. A large appeal of mechanistic interpretability is that it provides causal explanations for models’ outputs, i.e., a decision or prediction can be attributed in a causal manner to some component of the input. For example, causal tracing, a prominent tool in this line of work, perturbs internal activations in a manner that allows us to determine whether they cause a change in the output. This allows researchers to move beyond correlation-based explanations and instead understand the actual computational mechanisms driving observed model behavior, which in turn enables more precise debugging, bias detection, and control over generative outputs. Mechanistic interpretability research has uncovered a number of interesting and useful properties of DGMs. For example, specific neurons or layers in GANs and VAEs encode “disentangled” (i.e., distinct and interpretable) features, such as shape, texture, or pose in images (Shen et al., 2020; Mita et al., 2021). Other works have found that certain attention heads in transformers correspond to meaningful linguistic patterns, e.g., some might focus on syntactic structure while others might capture semantic information (Vig & Belinkov, 2019; Elhage et al., 2021). Such properties not only help practitioners better understand DGMs, they also enable them to control generation to some extent (Härkönen et al., 2020; Fetty et al., 2020).
@@
However, the reliability and comprehensiveness of mechanistic interpretability methods remain a subject of debate (Golechha & Dao, 2024a; Sharkey et al., 2025). A central criticism is that, while these techniques aim to identify causal relationships between model components and outputs, they may only provide partial or even misleading insights into the actual computational processes at work. For instance, attention patterns—which have been used by various methods to attribute model predictions to certain tokens (Xu et al., 2015; Choi et al., 2016, inter alia)—do not always faithfully reflect how or why certain tokens influence the final prediction (Jain & Wallace, 2019; Liu et al., 2022b). In some cases, they may merely highlight correlations rather than reveal deeper causal structures. Similar concerns have been raised about other methods for explaining model behaviors from mechanistic interpretability. Further, the new wave of large-scale generative models makes the application of some of these methods more difficult: the larger a model becomes, the (arguably) more difficult it becomes to fully reverse-engineer a prediction, both from a theoretical and computational standpoint. Every nuance of these models’ decision-making may not be deducible from a subset of neurons, layers, or attention heads, and interpretations derived from one subset of model components may overlook equally critical interactions elsewhere in the large network.
Representation engineering (Zou et al., 2023a) and the use of sparse autoencoders (Bricken et al., 2023) are lines of research in mechanistic interpretability that potentially address the former issues, offering interpretable explanations even for large-scale models. Efficient methods for circuit identification have started to address the latter issue (Hsu et al., 2025). However, we are still in need of validation methods that can confirm whether the identified “mechanisms” truly govern a model’s outputs, or whether they merely reflect convenient, yet incomplete, narratives about its internal workings.